Elastic collission in two dimenshan,write expression
1
Expert's answer
2012-12-04T08:42:27-0500
Let two objects with mass m1 and m2.
Initial velocity of object 1 = u1,
Initial velocity of object 2 = u2
Let v1 and v2 be the final velocities after elasticcollision, respectively.
Applying conservation of momentum principle, we get:
m1 * u1 + m2 * u2 = m1 * v1 + m2 * v2
Also the principle of kinetic energy conservation couldbe applied:
m1 * u1^2 / 2 + m2 * u2^2 / 2 = m1 * v1^2 / 2 + m2 * v2^2/ 2.